OpenStudy (juscallmesteve):

Differentiate y=(x+8)^sinx the equation so it is easier to read in post. I will also post my steps and would like someone to tell me if I am correct. (many steps will take some time)

OpenStudy (juscallmesteve):

\[y=(x+8)^{sinx}\]

OpenStudy (juscallmesteve):

step 1: \[y=(x+8)^{sinx}*ln(x+8)(sinx)\]

zepdrix (zepdrix):

Woah, I don't understand that first step...

zepdrix (zepdrix):

It looks like you tried to apply.... power rule... and exponential rule at the same time or something? Hmm

OpenStudy (juscallmesteve):

we were taught something like \[U(X)^{v(x)} = U(X)^{V(X)}*(\ln(U(X))*V(X))\]

OpenStudy (juscallmesteve):

for that step 1

zepdrix (zepdrix):

Hmm

OpenStudy (daniel.ohearn1):

Right. So first step would be take the natural log of both sides.

OpenStudy (juscallmesteve):

I have many other steps, but no need to go further if they are all wrong

OpenStudy (juscallmesteve):

so ln(y) = ln(x+8)^sinx

zepdrix (zepdrix):

\[\large\rm y=u^v\qquad\to\qquad \ln y=\ln(u^v)\qquad\to\qquad \ln y=v \ln(u)\]Yah you can do logarithms to help you along.

OpenStudy (daniel.ohearn1):

\[\ln y = \ln(x+8)\sin(x)\]

OpenStudy (daniel.ohearn1):

So..

OpenStudy (juscallmesteve):

is that * sinx? or

OpenStudy (daniel.ohearn1):

Yes

OpenStudy (juscallmesteve):

so \[\frac{ 1 }{ y }y \prime = \frac{ 1 }{ x+8 } * sinx\]

OpenStudy (daniel.ohearn1):

You got to apply the full on product rule

OpenStudy (juscallmesteve):

ok but why would you do the product rule before doing what ever the Ln rule is called, where the derivative of ln is 1/x * d/dx of x

OpenStudy (daniel.ohearn1):

You have two functions on the right side of the equation you have a product of, so you did half of it. The deriv of the ln part times the sinx ...

OpenStudy (juscallmesteve):

ok I will post my results shortly

OpenStudy (juscallmesteve):

\[\ln(y) = \ln(x+8)(cosx) + (\frac{ 1 }{ x+8 }*(x+8)\prime)(sinx)\]

OpenStudy (juscallmesteve):

\[\frac{ 1 }{ y }y \prime = \ln(x+8)(cosx) + (\frac{ 1 }{ x+8 })*sinx\]

OpenStudy (juscallmesteve):

that would reduce to \[\frac{ 1 }{ y }y \prime = \ln(x+8)(cosx)+(\frac{ sinx }{ cosx })\]

OpenStudy (juscallmesteve):

messed up on that last one

OpenStudy (daniel.ohearn1):

When you differentiate you are trying to find the derivative y prime, right? So all you need to do is multiply each side by y...

OpenStudy (juscallmesteve):

\[\frac{ 1 }{ y }y \prime = \ln(x+8)(cosx) + (\frac{ sinx }{ x+8 })\]

OpenStudy (daniel.ohearn1):

Keeping the function in terms of x

OpenStudy (juscallmesteve):

not sure what you mean

OpenStudy (daniel.ohearn1):

so ideally you end up having y'=...... noting that y=f(x) was already defined at the beginning of the problem

OpenStudy (juscallmesteve):

yes that is what I want lol

OpenStudy (bobo-i-bo):

This identity gives a simple way to do the differential :P \[a^b=(e^{\ln(a)})^b=e^{bln(a)}\]

OpenStudy (juscallmesteve):

so then \[e ^{sinx \ln(x+8)}\]

OpenStudy (bobo-i-bo):

yup :) Then apply chain rule

OpenStudy (juscallmesteve):

to just the exponent? or do exponent rule and then chain?

OpenStudy (bobo-i-bo):

So use the chain rule after using my identity. You've got to try and think of a good substitution to use :)

OpenStudy (juscallmesteve):

then would it be e\[e ^{\frac{ sinx }{ x+8 }+(cosx)(\ln(x+8))}\]

OpenStudy (bobo-i-bo):

Nope

OpenStudy (bobo-i-bo):

\[y=(x+8)^{sinx}=e^{\sin(x)\ln(x+8)}\]

OpenStudy (bobo-i-bo):

Use the substitution \[u=\sin(x)\ln(x+8)\] for the chain rule

OpenStudy (juscallmesteve):

then u = \[\frac{ sinx }{ x+8 } + (cosx)(\ln(x+8))\]

OpenStudy (bobo-i-bo):

$$\frac{du}{dx}$$ is equal to that, yes. Not u.

OpenStudy (juscallmesteve):

i used product rule

OpenStudy (juscallmesteve):

I am sorry, but I am just completely lost at where we are.

OpenStudy (juscallmesteve):

I really do appreciate the help, I will go watch some youtube videos and then come back

OpenStudy (bobo-i-bo):

We wish to find \[\frac{dy}{dx}\]. But \[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\] by the chain rule. Now, if we let \[u=sin(x)ln(x+8)\] then: \[y=e^u\] since \[y=e^{sin(x)ln(x+8)}\] So we can find \[\frac{dy}{dx}\] by finding \[\frac{dy}{du}\] and \[\frac{du}{dx}\]

OpenStudy (juscallmesteve):

Thank you for that information

OpenStudy (bobo-i-bo):

:)